OpenStudy (anonymous):

Solve: (x + 5)^2 = 16 A. -1, -9 B. -1, 9 C. 1, -9 D. 1, 9

OpenStudy (anonymous):

If you multiply (x+5) by (x+5) then you get x^2+10x+25 = 16 Set it equal to 0: 0 = x^2+10x+9 This will factor down to: (x+9)(x+1) Which makes the solutions: x=-9 and x=-1